# Data-Driven Exploration for a Class of Continuous-Time Linear–Quadratic Reinforcement Learning Problems

### v1

@@ Abstract
We study reinforcement learning (RL) for the same class of continuous-time stochastic linear–quadratic (LQ) control problems as in [1], where volatilities depend on both states and controls while states are scalar-valued and running control rewards are absent. We propose a model-free, data-driven exploration mechanism that adaptively adjusts entropy regularization by the critic and policy variance by the actor. Unlike the constant or deterministic exploration schedules employed in [1], which require extensive tuning for implementations and ignore learning progresses during iterations, our adaptive exploratory approach boosts learning efficiency with minimal tuning. Despite its flexibility, our method achieves a sublinear regret bound that matches the best-known model-free results for this class of LQ problems, which were previously derived only with fixed exploration schedules. Numerical experiments demonstrate that adaptive explorations accelerate convergence and improve regret performance compared to the non-adaptive model-free and model-based counterparts.
Keywords: C ontinuous-Time Reinforcement Learning, Stochastic Linear–-Quadratic Control, Actor, Critic, Data-Driven Exploration, Model-Free Policy Gradient, Regret Bounds.
@@ 1 Introduction
Linear–quadratic (LQ) control is a cornerstone of optimal control theory, widely applied in engineering, economics, and robotics due to its analytical tractability and practical relevance. Its theory has been extensively developed in the classical model-based setting where all the LQ coefficients are known and given [2, 3]. However, in real-world applications, complete knowledge of model parameters is rarely available. Many environments exhibit LQ-like structure yet key parameters in dynamics and objectives may be partially known or entirely unknown. While one can use historical data to estimate those model parameters as commonly done in the so-called “plug-in” approach, estimating some of the parameters to a required accuracy may not be possible. For instance, it is statistically impossible to estimate a stock return rate for a very small period – a phenomenon termed “mean-blur” [4, 5]. On the other hand, objective function values can be highly sensitive to estimated model parameters, rendering inferior performances or even complete irrelevance of model-based solutions [6].
@@
Given these challenges, model-free reinforcement learning (RL) has emerged as a promising alternative for stochastic control, which directly learns optimal policies through data without estimating model parameters. A centerpiece of RL is exploration, with which the agent interacts with the unknown environment and improves her policies. The simplest exploration strategy is perhaps the so-called $\epsilon$-greedy for bandits problem [7]. More sophisticated and efficient exploration strategies include intrinsic motivation methods that reward exploring novel or unpredictable states [8, 9, 10, 11, 12, 13, 14], count-based exploration that encourages the agent to explore less frequently visited states [15, 16], go-explore that explicitly stores and revisits promising past states to enhance sample efficiency [17, 18], imitation-based exploration which leverages expert demonstrations [19], and safe exploration methods that rely on predefined safety constraints or risk models [20, 21]. While these methods are shown to be successful in their respective contexts, they are all for discrete-time problems and seem difficult to be extended to the continuous-time counterparts. However, most real-life applications are inherently continuous-time with continuous state spaces and possibly continuous control spaces (e.g., autonomous driving, robot navigation, and video game playing). On the other hand, transforming a continuous-time problem into a discrete-time one upfront by discretization has shown to be problematic when the time step size becomes very small [22, 23, 24].
@@
Thus, there is a pressing need for developing exploration strategies tailored to continuous-time RL, particularly and foremost in the LQ setting, for achieving both stability and learning efficiency. Recent works, such as [1, 25, 26], have employed constant or deterministic exploration schedules to regulate exploration parameters and proved to achieve sublinear regret bounds. However, these approaches come with notable drawbacks. They require extensive manual tuning of the exploration hyperparameters, which considerably increases computational costs, yet these tuning efforts are typically not reflected in theoretical analyses and results including those concerning regret bounds. Moreover, pre-determined exploration strategies lack adaptability by nature, as they do not react to the incoming data nor to the current learning states, typically resulting in slower convergence and increased variance from excessive exploration or premature convergence to suboptimal policies from insufficient exploration.
@@
This paper presents a companion research of [1]. We continue to study the same class of stochastic LQ problems in [1], but develop a data-driven adaptive exploration framework for both critic (the value functions) and actor (the control policies). Specifically, we propose a data-driven adaptive exploration strategy that dynamically adjusts the entropy regularization parameter for the critic and the stochastic policy variance for the actor based on the agent’s learning progress. Moreover, we provide theoretical guarantees by proving the almost sure convergence of both the adaptive exploration and policy parameters, along with a sublinear regret bound of $O(N^{\frac{3}{4}})$. This bound matches that in [1] whose analysis relies on fixed/deterministic exploration schedules. Finally, we validate our theoretical results through numerical experiments, demonstrating a faster convergence and improved regret bounds compared with a model-based benchmark with fixed exploration. For a comparison with the previous paper [1], we design experiments with both fixed and randomly generated model parameters. While theoretically the regret bounds in the two papers are of the same order, the numerical experiments show that our adaptive exploration strategy consistently outperforms and achieves lower regrets.
@@ 2.1 Stochastic LQ Control: Classical Setting
where $Q\geq 0$ and $H\geq 0$ are scalar weighting parameters. If the model parameters $A$, $B$, $C_{j}$, $D_{j}$, $Q$, and $H$ are known, the problem has an explicit solution as in [3, Chapter 6].
@@ 3.2 Adaptive Critic Exploration
In the related literature, $\gamma$ has been either taken as an exogenous constant hyperparameter ([1, 28, 29]) or a deterministic sequence of $n$ ([25]), which in turn requires extensive tunings in implementations. By contrast, we derive an adaptive temperature schedule, which is endogenous and data-driven, adjusted dynamically over iterations. Specifically, the adaptive update of $\gamma_{n}$ is determined by the values of the random processes $\boldsymbol{\theta}_{n}$ and a predefined deterministic sequence $b_{n}$ as follows
@@ 5.1.3 Almost Sure Convergence of $\gamma_{n}$ and $\Gamma_{n}$
By the assumptions (i)-(ii) of (38), as well as (41) and (42), we know $\sum\kappa^{(\Gamma)}_{n}<\infty$ and $\sum\eta^{(\Gamma)}_{n}<\infty$. It then follows from [38, Theorem 1] that $\left|U^{(\Gamma)}_{n}\right|^{2}$ converges to a finite limit and $\sum\zeta^{(\Gamma)}_{n}<\infty$ almost surely.
@@ 5.1.4 Convergence Rate of $\Gamma_{n}$
$\mathbb{E}[|\Gamma_{n+1}|^{2}]\leq c\frac{(e\vee\log n)^{p_{2}}(1\vee\log\log n)^{\frac{4}{3}}}{n^{\frac{1}{2}}}$
@@
$\mathbb{E}\left[|U^{(\Gamma)}_{n+1}|^{2}\Big{|}\mathcal{G}_{n}\right]\leq(1-\tilde{c}\hat{a}_{n})|U^{(\Gamma)}_{n}|^{2}+4\hat{a}_{n}^{2}\hat{\eta}^{(\Gamma)}_{n}$
, where
@@
where $c$ and $p_{2}$ are positive constants. Thus it follows that
$\mathbb{E}[|\Gamma_{n+1}-\Gamma_{n}^{*}|^{2}]\leq c\frac{(e\vee\log n)^{p_{2}}(1\vee\log\log n)^{\frac{4}{3}}}{n^{\frac{1}{2}}}.$
@@ 7 Conclusions
This paper is a continuation of [1], aiming to develop a model-free continuous-time RL framework for a class of stochastic LQ problems with state- and control-dependent volatilities. A key contribution, compared to [1], is the introduction of a data-driven exploration mechanism that adaptively adjusts both the temperature parameter controlled by the critic and the stochastic policy variance managed by the actor. This approach overcomes the limitations of fixed exploration schedules, which often require extensive tuning and incur unnecessary exploration costs. While the theoretically proved sublinear regret bound of $O(N^{\frac{3}{4}})$ is the same as that in [1], numerical experiments confirm that our adaptive exploration strategy significantly improves learning efficiency compared to [1].
@@ Appendix A More Details for Numerical Experiments
For all the plots in the first and second experiments, for both our adaptive model-free algorithm and model-based benchmarks, regression lines are fitted using data from iterations 5,000 to 100,000. This avoids the influence of early-stage learning noise. In all the regret plots, we use the median and exclude extreme values. [^5] Similar conclusions hold when using the mean instead.

### v2

# Data-Driven Exploration for a Class of Continuous-Time Indefinite Linear–Quadratic Reinforcement Learning Problems

@@ Abstract
We study reinforcement learning (RL) for the same class of continuous-time stochastic linear–quadratic (LQ) control problems as in [1], where volatilities depend on both states and controls while states are scalar-valued and there are no running control rewards, the last feature also leading to the so-called indefinite LQ controls. We propose a model-free, data-driven exploration mechanism that adaptively adjusts entropy regularization by the critic and policy variance by the actor. Unlike the constant or deterministic exploration schedules employed in [1], which require extensive tuning for implementations and ignore learning progresses during iterations, our adaptive exploratory approach boosts learning efficiency with minimal tuning. Despite its flexibility, our method achieves a sublinear regret bound that matches the best-known model-free results for this class of LQ problems, which were previously derived only with fixed exploration schedules. Numerical experiments demonstrate that adaptive explorations accelerate convergence and improve regret performance compared to the non-adaptive model-free and model-based counterparts.
Keywords: S tochastic Indefinite Linear–-Quadratic Control, Continuous-Time Reinforcement Learning, Actor, Critic, Data-Driven Exploration, Model-Free Policy Gradient, Regret Bounds.
@@ 1 Introduction
Linear–quadratic (LQ) control is a cornerstone of optimal control theory, widely applied in engineering, economics, and robotics among many others due to its analytical tractability and practical relevance. Its theory has been extensively developed in the classical model-based setting where all the LQ coefficients are known and given [2, 3]. Assuming the objective is to minimize the cost functional, a key assumption in the LQ theory is that the control cost weighting matrix must be positive definite. Intuitively, this is because if increasing the level of control is costless or even beneficial, then the optimal control would just be infinitely large leading to an ill-posed problem. [4] finds, however, that certain stochastic control LQ control problems with indefinite cost matrices—including those associated with control—can still be well-posed if the diffusion coefficients depend on the control variable. This seemingly surprising result actually has a simple explanation: while benefiting from an indefinite control cost, a larger control is disadvantageous in terms of increasing the level of uncertainty because the volatility term depends on control multiplicatively. Therefore, one needs to carefully choose the optimal level of control, rendering a well-posed optimization problem. [4] coins the term “indefinite stochastic LQ control” for such a problem, and establishes conditions for well-posedness and solvability via newly introduced generalized Riccati equations. Subsequent works extend the theory in several directions, including infinite-time horizons [5] and connections with linear matrix inequalities [6], asymptotic properties of the Riccati solutions [7], and semidefinite programming for numerical computations [8]. Research on indefinite LQ controls and the related generalized Riccati equations has been active to this day; see, e.g., [9, 10, 11, 12].
@@
All the research on indefinite LQ controls, and indeed most studies on stochastic controls, are undertaken in a model-based paradigm where the system coefficients and objective/cost functionals are fully known. In real-world applications, however, complete knowledge of model parameters is rarely available. Many environments exhibit LQ-like structure yet key parameters in dynamics and objectives may be partially known or entirely unknown. While one can use historical data to estimate those model parameters as commonly done in the so-called “plug-in” approach, estimating some of the parameters to a required accuracy may not be possible. For instance, it is statistically impossible to estimate a stock return rate for a very small period – a phenomenon termed “mean-blur” [13, 14]. On the other hand, objective function values can be highly sensitive to estimated model parameters, rendering inferior performances or even complete irrelevance of model-based solutions [15].
@@
Given the limitations of model-based controls, particularly the challenges associated with unknown model coefficients, reinforcement learning (RL) rises to provide a promising remedy. Simply put, RL is stochastic control with unknown model parameters. Yet it never attempts to estimate any model coefficients, and solves the problem in a fundamentally different way compared to the classical, model-based stochastic control. A centerpiece of RL is exploration, with which the agent interacts with the unknown environment and improves her policies. The simplest exploration strategy is perhaps the so-called $\epsilon$-greedy for bandits problem [16]. More sophisticated and efficient exploration strategies include intrinsic motivation methods that reward exploring novel or unpredictable states [17, 18, 19, 20, 21, 22, 23], count-based exploration that encourages the agent to explore less frequently visited states [24, 25], go-explore that explicitly stores and revisits promising past states to enhance sample efficiency [26, 27], imitation-based exploration which leverages expert demonstrations [28], and safe exploration methods that rely on predefined safety constraints or risk models [29, 30]. While these methods are shown to be successful in their respective contexts, they are all for discrete-time problems and seem difficult to be extended to the continuous-time counterparts. However, most real-life applications are inherently continuous-time with continuous state spaces and possibly continuous control spaces (e.g., autonomous driving, robot navigation, and video game playing). On the other hand, transforming a continuous-time problem into a discrete-time one upfront by discretization has shown to be problematic when the time step size becomes very small [31, 32, 33].
@@
Thus, there is a pressing need for developing exploration strategies tailored to continuous-time RL, particularly and foremost in the LQ setting, for achieving both stability and learning efficiency. Recent works, such as [1, 34], have employed constant or deterministic exploration schedules to regulate exploration parameters and proved to achieve sublinear regret bounds. However, these approaches come with notable drawbacks. They require extensive manual tuning of the exploration hyperparameters, which considerably increases computational costs, yet these tuning efforts are typically not reflected in theoretical analyses and results including those concerning regret bounds. Moreover, pre-determined exploration strategies lack adaptability by nature, as they do not react to the incoming data nor to the current learning states, typically resulting in slower convergence and increased variance from excessive exploration or premature convergence to suboptimal policies from insufficient exploration.
@@
This paper presents a companion research of [1]. We continue to study the same class of stochastic LQ problems in [1], which is also of an indefinite LQ type because control is absent in the objective functional. We develop a data-driven adaptive exploration framework for both critic (the value functions) and actor (the control policies). Specifically, we propose an adaptive exploration strategy that dynamically adjusts the entropy regularization parameter for the critic and the stochastic policy variance for the actor based on the agent’s learning progress. Moreover, we provide theoretical guarantees by proving the almost sure convergence of both the adaptive exploration and policy parameters, along with a sublinear regret bound of $O(N^{\frac{3}{4}})$. This bound matches the best-known model-free result from [1], which assumes a nonzero initial state and uses a fixed/deterministic exploration schedule, whereas we removes this assumption and learns entirely from data. Finally, we validate our theoretical results through numerical experiments, demonstrating a faster convergence and improved regret bounds compared with a model-based benchmark with fixed exploration. For a comparison with the previous paper [1], we design experiments with both fixed and randomly generated model parameters. While theoretically the regret bounds in the two papers are of the same order, the numerical experiments show that our adaptive exploration strategy consistently outperforms and achieves lower regrets.
A distinctive advantage of the RL approach over the traditional model-based one for indefinite LQ control is that the latter relies heavily on the generalized (or indefinite) Riccati equation, which is a highly complex and unconventional Riccati equation. In particular, it involves a positive definiteness constraint and may be singular due to the indefiniteness of the control weighting cost [4]. A central task for solving indefinite LQ control problems has been to solve this equation analytically or numerically (e.g. [6, 7, 8, 35, 36]), which remains open for the most general case in terms of its well-posedness. However, RL does not need to involve this equation at all, in the same way that RL gets around the HJB PDE for a general stochastic control problem.
@@ 2.1 Stochastic LQ Control: Classical Setting
where $Q\geq 0$ and $H\geq 0$ are scalar weighting parameters. Note that here we maximize the reward functional (but with the negative signs) – instead of minimizing a cost functional as in the usual LQ literature – to be consistent with the setting of [1]. If the model parameters $A$, $B$, $C_{j}$, $D_{j}$, $Q$, and $H$ are known, the problem has an explicit solution as in [3, Chapter 6].
@@ 3.2 Adaptive Critic Exploration
In the related literature, $\gamma$ has been either taken as an exogenous constant hyperparameter ([38, 1]) or a deterministic sequence of $n$ ([34]), which in turn requires extensive tunings in implementations. By contrast, we derive an adaptive temperature schedule, which is endogenous and data-driven, adjusted dynamically over iterations. Specifically, the adaptive update of $\gamma_{n}$ is determined by the values of the random processes $\boldsymbol{\theta}_{n}$ and a predefined deterministic sequence $b_{n}$ as follows
@@ 5.1.3 Almost Sure Convergence of $\gamma_{n}$ and $\Gamma_{n}$
By the assumptions (i)-(ii) of (38), as well as (41) and (42), we know $\sum\kappa^{(\Gamma)}_{n}<\infty$ and $\sum\eta^{(\Gamma)}_{n}<\infty$. It then follows from [47, Theorem 1] that $\left|U^{(\Gamma)}_{n}\right|^{2}$ converges to a finite limit and $\sum\zeta^{(\Gamma)}_{n}<\infty$ almost surely.
@@ 5.1.4 Convergence Rate of $\Gamma_{n}$
$\mathbb{E}[|\Gamma_{n}|^{2}]\leq c\frac{(e\vee\log n)^{p_{2}}(1\vee\log\log n)^{\frac{4}{3}}}{n^{\frac{1}{2}}}$
@@
$\mathbb{E}\left[|U^{(\Gamma)}_{n+1}|^{2}\Big{|}\mathcal{G}_{n}\right]\leq(1-\tilde{c}\hat{a}_{n})|U^{(\Gamma)}_{n}|^{2}+4\hat{a}_{n}^{2}\hat{\eta}^{(\Gamma)}_{n},$
where
@@
where $c$ and $p_{2}$ are positive constants. Since $n\geq 1$, it follows that
$\mathbb{E}[|\Gamma_{n+1}-\Gamma_{n+1}^{*}|^{2}]\leq c\frac{(e\vee\log n)^{p_{2}}(1\vee\log\log n)^{\frac{4}{3}}}{n^{\frac{1}{2}}}$
$\leq$ $c\frac{(e\vee\log(n+1))^{p_{2}}(1\vee\log\log(n+1))^{\frac{4}{3}}}{(n+1)^{\frac{1}{2}}}\frac{(n+1)^{\frac{1}{2}}}{n^{\frac{1}{2}}}$
$\leq$ $c\sqrt{2}\frac{(e\vee\log(n+1))^{p_{2}}(1\vee\log\log(n+1))^{\frac{4}{3}}}{(n+1)^{\frac{1}{2}}}.$
@@ 7 Conclusions
This paper is a continuation of [1], aiming to develop a model-free continuous-time RL framework for a class of indefinite stochastic LQ problems with state- and control-dependent volatilities. A key contribution, compared to [1], is the introduction of a data-driven exploration mechanism that adaptively adjusts both the temperature parameter controlled by the critic and the stochastic policy variance managed by the actor. This approach overcomes the limitations of fixed exploration schedules, which often require extensive tuning and incur unnecessary exploration costs. While the theoretically proved sublinear regret bound of $O(N^{\frac{3}{4}})$ is the same as that in [1], numerical experiments confirm that our adaptive exploration strategy significantly improves learning efficiency compared to [1].
@@ Appendix A More Details for Numerical Experiments
For all the plots in the first and second experiments, for both our adaptive model-free algorithm and model-based benchmarks, regression lines are fitted using data from iterations 5,000 to 100,000. This avoids the influence of early-stage learning noise. In all the regret plots, we use the median and exclude extreme values; similar conclusions hold when using the mean instead.